Presentation is loading. Please wait.

Presentation is loading. Please wait.

MOM Report Paul Soler MICE Operations Manager University of Glasgow MICO April 20, 2009.

Similar presentations

Presentation on theme: "MOM Report Paul Soler MICE Operations Manager University of Glasgow MICO April 20, 2009."— Presentation transcript:

1 MOM Report Paul Soler MICE Operations Manager University of Glasgow MICO April 20, 2009

2 Commissioning Needs-Soler2 Commissioning Needs Working Decay Solenoid – Stands.. Many thanks to Matt, Jim, Mike, Willie, Linda, …. —Cooldown completed Thursday 16 April, temperature ~4.5K —Matt attempted turning on power supply on Friday 17 April (14:30) but quench protector light number 5 tripped o Helium vented from cryostat relief back up to buffer tank as expected (fridge trip signal not disabled). o Dewar emptied and temperature initially raised to 10.5-11.5K but by 17:00 dewar refilled to 80% and temperature back to 4.5K —One vacuum pump tripped over weekend (but isolated from circuit) —Successfully turned on power supply and ramped up to 400 A today o Evidence that all coils have gone superconducting (0V at stable current) o First indication that solenoid fix (ie. fitting Multi-Layer Installation) has solved quenching problem in magnet o Quench Protector keeps on tripping at turn on (it was bypassed for test) – light number 5 still trips o Solenoid warning lights now also working when AC contact on/off o Need to discuss when to go to full current (870 A) —Need PSI expert to ensure Quench Protector working

3 April 20, 2009Commissioning Needs-Soler3 Solenoid Control System

4 April 20, 2009Commissioning Needs-Soler4 DS Cooldown Sequence 4.6K

5 April 20, 2009Commissioning Needs-Soler5 DS Ramp Up 400 A* * Current badly calibrated in control system: 332A in plot corresponds to 400 A Polarity of circuit for C10 seems to be reversed Total voltage all coils

6 April 20, 2009Commissioning Needs-Soler6 Commissioning Needs Working target – Stands.. —one in R78 for demo —one in ISIS for data Detectors – —luminosity monitor o Brunel & Paul Soler working on this - ~£1300 for new PMTs —Tracker cryostats o Test Pressure Release valve by Alan this week – still to happen o Install emergency shutoff valves on cryostats – all 4 installed and working o Install exhaust filter – this week —CKOV o New temp and humidity monitoring device installed o Need network connection and power in the DSA – installed but info not yet logged (next week) o HV power supply crate died (spare) – removed and sent for repair Get Decay Solenoid magnet-on lights working —repositioned light on top of turret – both lights now working Air conditioning —6 holes in West Wall for A/C venting finished. Need supports for pipes

7 April 20, 2009Commissioning Needs-Soler7 Issues Main commissioning issue: when to remove shielding? —Need to make decision by 23 April —Willie suggests we wait till August shutdown to remove shielding since quench protection will not be verified Proposed schedule by Willie: 1.Wednesday 19th August erect scaffold 2.Twenty wooden pallets to be delivered to Synchrotron 3.Thursday 20th August commence removal of shield steel plate (Total mass 22.4 Tonne) 4.From Thursday - Sunday we need to remove 5.6 tonne's per/day 5.Monday 24th August strip down support frame 6.Monday 24th August re-position Dipole 1 7.Monday 24th August make a request to ISIS Teams Mechanical & Vacuum for the coupling of Water works to Dipole 1 and the re-installation of vacuum pipe work and Ion pump to commence 25th August Can accommodate schedule also if we have to wait 1 week for radiation cooldown Parallel work on target should not be issue since not in same area, different people involved and crane only needed for 1 day for target installation

8 April 20, 2009Commissioning Needs-Soler8 Issues Documentation inventory/overhaul - ongoing —Instructions for shifters – old documentation needs updating —Safety – need all relevant documents in one obvious place Computing —Need network connections to progress in Hall —Still cannot print on OKI printer in R76 —Wireless in R76 still not installed but network people aware of problem —Need Windows PC in MLCR – still pending —Need connection to GRID from MLCR – pending Need beamline magnets on to test alarm handler —Water will be restored to magnets – during this week

Download ppt "MOM Report Paul Soler MICE Operations Manager University of Glasgow MICO April 20, 2009."

Similar presentations

Ads by Google